### v1

@@ 7 Experiments
We have conducted extensive experiments on synthetic and real images to evaluate the performance of our proposed Quasiconformal Interactive Segmentation model. In this section, we will report the experimental results. All implementations were executed using Matlab R2022b on a Windows 11 x64 platform with a 3.20 GHz AMD Ryzen 5800H processor and 16 GB RAM. In our experiments, all images were resized to $512\times 512$ and their intensities were rescaled to $[0,255]$. The weight for the Laplacian term is $\alpha_{1}=0.001$, and the weight for the regularization of the Beltrami coefficients is $\alpha_{2}=100$.
@@ 7.1 Interactive segmentation of synthesized images
We first test our proposed model on simple synthetic images. Figure 4 illustrates the quasiconformal interactive segmentation of an image featuring a taco on a plate. Our goal is to segment the taco. The initial contour is a simple rectangle, as shown in (a). The initial contour defines the desired topology of the segmentation mask, which is simply-connected. The initial segmentation result using the quasiconformal segmentation model is displayed in (b), which includes the plate as well. The deformation $\phi$ that transforms the initial contour to the initial result is shown in (c). To exclude the plate from the segmentation mask, we introduce a negative click on the plate, represented by the red dot in (b). The click map associated with this click is shown in (d). The quasiconformal interactive segmentation result is presented in (e), accurately segmenting the desired taco. The deformation corresponding to the final interactive segmentation result is shown in (f), which is bijective. Thus, the segmentation mask preserves the same topology as the initial contour..
Figure 5 presents another synthetic image featuring two eating utensils: a fork and a knife. Our objective is to segment the knife. (a) shows the initial contour, which is a simple rectangle. Again, the initial contour defines the desired topology of the segmentation mask, which is simply-connected. The initial segmentation result using the quasiconformal segmentation model is displayed in (b), with its corresponding deformation shown in (c). Clearly, the segmentation mask includes some undesired regions. To exclude these undesired regions, two negative clicks are introduced, indicated by the red dots in (b). The click map associated with these clicks is shown in (d). The quasiconformal interactive segmentation result is presented in (e), accurately segmenting the desired knife. The deformation corresponding to the interactive segmentation result is shown in (f), which is bijective. Thus, the segmentation mask preserves the same topology as the initial contour.
@@ 7.2 Interactive segmentation of medical images
In this subsection, we evaluate our proposed segmentation model on real medical images. More specifically, we use brain magnetic resonance images (MRIs) from the BraTs21 dataset in our experiment.
@@
Figure 6 shows an MRI of a brain with a lesion region. Our objective is to segment the lesion region. (a) presents the initial contour, which is a simply-connected polygon. (b) depicts the initial segmentation mask obtained using the quasiconformal segmentation model, and the corresponding deformation is shown in (c). Clearly, the segmentation mask contains some undesired regions that do not belong to the lesion region. To exclude these undesired regions, two negative clicks are introduced and indicated by the red dots in (b). (d) shows the corresponding click map. By employing our proposed QIS model, the segmentation result is presented in (e), accurately segmenting the lesion. The corresponding deformation is shown in (f), which is bijective. Thus, the segmentation mask preserves the same topology as the initial contour.
Figure 7 presents another example of interactively segmenting the lesion region of a brain MRI. The initial segmentation result with a simply-connected initial contour using the quasiconformal segmentation model is displayed in (b), and the corresponding deformation is shown in (c). Once again, the initial segmentation mask contains some undesired regions that do not belong to the lesion region. Two negative clicks, indicated as red dots in (b), are introduced to exclude these undesired regions. The click map is shown in (d). By incorporating the click map, the QIS model accurately segments the lesion region. The segmentation mask preserves the same topology as the initial contour, which is also evident from the bijective deformation shown in (f).
@@
Figure 8 and Figure 9 show the results of QIS on two noisy brain MRIs, with the objective of extracting the lesion region. In each figure, (a) displays the initial contour, while the initial segmentation result using the quasiconformal segmentation model is shown in (b). The corresponding segmentation mask is displayed in (c). In both cases, the initial segmentation mask contains some undesired regions. To exclude these undesired regions, negative clicks are introduced, indicated by the red dots in (b). By incorporating the corresponding click map in (d), our proposed QIS model accurately segments the lesion region. The segmentation mask preserves the same topology as the initial contour, as evidenced by the bijective deformation shown in (f). These examples illustrate the effectiveness of our proposed QIS model, even on noisy images.
@@ 7.3 Comprehensive analysis of QIS model
Comparison with modified Chan-Vese segmentation model
@@
The QIS model is a deformation-based segmentation model incorporating the quasiconformal segmentation model with the click map. One advantage of the quasiconformal segmentation model is that it ensures the segmentation mask maintains the prescribed topology, avoiding topological errors and enhancing the accuracy and efficiency of the interactive segmentation model. To investigate the importance of incorporating the quasiconformal segmentation model, we compare our method with an alternative method that integrates the level-set segmentation model with the click map. Instead of utilizing the quasiconformal segmentation model, we can run the Chan-Vese segmentation model with $I+rM_{clicks}$ as the input image using a suitable parameter $r$. The segmentation results of several natural images using the modified Chan-Vese model are presented in Figure 10. The first row of Figure 10 illustrates the initial segmentation outcomes of four natural images using the conventional Chan-Vese segmentation model. The two images on the left are clean, while the two on the right are noisy. The second row displays the initial segmentation results of the same four natural images using the quasiconformal segmentation model. In both methods, the initial segmentation masks include some undesired regions and exclude certain regions of interest. Some positive clicks (in blue) and negative clicks (in red) are introduced to guide the segmentation. The third row shows the segmentation results obtained by the modified Chan-Vese segmentation method, incorporating click maps. The segmentation masks exhibit improved accuracy in segmenting the regions of interest, though they are not entirely precise. Finally, the last row demonstrates the segmentation results obtained by the QIS model, which accurately segments the regions of interest. The topology of the segmentation mask is also consistent with that of the initial contour. Note that the clicks inputted by users are the same for both the modified Chan-Vese segmentation model and the QIS model.
@@
Figure 11 presents another comparison using brain MRIs. The first row of Figure 11 shows the initial segmentation results of four brain MRIs using the conventional Chan-Vese segmentation model. Again, the two images on the left are clean, while the two on the right are noisy. Evidently, the initial segmentation masks contain some undesired regions and exclude some regions of interest. The second row shows the segmentation results obtained by the modified Chan-Vese segmentation method, incorporating click maps. The clicks inputted by the users are the same as those used by the QIS model, as shown in Figure 6, Figure 7, Figure 8 and Figure 9. The segmentation masks can better segment the lesion, although they are still not accurate. The last row shows the segmentation results obtained by the QIS model, which accurately segments the lesions. The topology of the segmentation mask is consistent with that of the initial contour.
@@
Figure 10: Segmentation results on natural images obtained using the QIS model and the modified Chan-Vese segmentation model with the integration of click maps. The left side displays the segmentation results on two clean images, while the right side shows the segmentation results on two noisy images.
@@
Figure 11: Segmentation results on brain MRIs obtained using the QIS model and the modified Chan-Vese segmentation model with the integration of the click map. The left side displays the segmentation results on two clean brain MRIs, while the right side shows the segmentation results on two noisy brain MRIs.
Comparison with GrabCut
@@
We compare the QIS model with a popular interactive segmentation model called GrabCut. Similar to the QIS model, GrabCut requires the user to provide interactive inputs of foreground and background brushes. The comparison between the QIS and GrabCut methods on a noisy image capturing flowers is shown in Figure 12. In Figure 12(a), the iterative process of the QIS model is displayed. The negative clicks in each iteration, represented by red brushes, are also shown. The QIS effectively segments the noisy image in four steps with only a few brushes introduced. Figure 12(b) illustrates the iterative process of GrabCut, where the brushes introduced by the user (red for foreground and blue for background) are also shown. GrabCut produces a reasonably accurate segmentation mask after seven steps, although it is evident that it is comparatively less accurate than QIS. Figure 13 shows a further comparison between GrabCut and QIS using another noisy image capturing a road scene, with the objective of segmenting the car. The iterative process of QIS is shown in (a), where the positive and negative clicks are denoted by red and blue brushes respectively. QIS accurately segments the car in just five steps. (b) displays the iterative process using GrabCut, with the segmentation result after eight steps exhibited, which is comparatively less accurate than QIS.
@@
Figure 13: Another comparison between the QIS and GrabCut segmentation methods. The top demonstrates the iterative segmentation process using the QIS method, with positive and negative clicks labeled as blue and red brushes respectively. The final segmentation result is achieved in 5 steps. The bottom depicts the iterative process using the GrabCut method, with positive and negative brushes labeled as blue and red respectively. The segmentation result after 8 steps is displayed.
@@ 8 Conclusion
In this paper, we propose the quasi-conformal interactive segmentation (QIS) model, with the incorporation of user’s interactions. User’s interactions are defined by two actions: a positive click and a negative click, represented by left and right clicks, respectively. These clicks serve to indicate the regions that the user wants to include or exclude from the segmentation process. Clicks are transformed into a click map, which is then integrated into the proposed model using a novel interactive segmentation energy term that we have designed. The segmentation mask is obtained by deforming a template mask with the same topology as the object of interest using an orientation-preserving quasiconformal mapping. This approach helps to avoid topological errors in the segmentation results. We provide a thorough analysis of the proposed model, including theoretical support for the ability of QIS to include or exclude regions of interest or disinterest based on the user’s indication. To evaluate the performance of QIS, we conduct experiments on synthesized images, medical images, natural images, and noisy natural images. The results demonstrate the efficacy of our proposed method.

### v2

@@ 6 Experiments
We have conducted extensive experiments on synthetic and real images to evaluate the performance of our proposed Quasi-conformal Interactive Segmentation (QIS) model. In this section, we will report the experimental results. All implementations were executed using Matlab R2022b on a Windows 11 x64 platform with a 3.20 GHz AMD Ryzen 5800H processor and 16 GB RAM. In our experiments, all images were resized into $512\times 512$ and their intensities were rescaled into $[0,255]$. The weight for the Laplacian term is $\alpha_{1}=0.001$, and the weight for the regularization of the Beltrami coefficients is $\alpha_{2}=100$. The implementation for the Chan-Vese model is from [44] with default parameters while the GrabCut is the built-in function of Matlab.
@@ 6.1 Interactive segmentation of synthesized images
We first evaluate our proposed model on simple synthetic images to provide a fundamental understanding of its functionality. To segment the taco in Figure 5, the model operates as follows. The user provides an initial contour (a) to approximate the region of interest and establish the desired topology of the segmentation mask. As described in Section 4.1, the initial segmentation (b) is obtained by deforming the initial contour using the mapping shown in (c). This initial segmentation includes both the taco and the plate. To exclude the plate, the user inputs a negative click (indicated by the red point in (b)). Using our proposed method for computing click maps, detailed in Section 4.2, we obtain the updated click map in (d). With this click map and the previous segmentation result, we solve the interactive segmentation energy term (4). The final result is refined and improved as shown in (e), using the quasi-conformal mapping illustrated in (f).
The second example, illustrated in Figure 6, involves segmenting two eating utensils: a fork and a knife. Following the same procedures, the initial contour, initial segmentation, and associated deformation mapping are provided in (a), (b), and (c), respectively. However, since the regions to be excluded are interrupted by the knife, two negative clicks are required because each click can only indicate one connected region by the click map algorithm used (Section 4.2). Therefore, the user must input two clicks to properly exclude both disconnected regions. The computed click map is presented in (d). The final segmentation result and its associated quasi-conformal mapping are shown in (e) and (f), respectively.
@@ 6.2 Interactive segmentation of medical images
In this subsection, we evaluate our proposed segmentation model on real medical images. More specifically, we use brain magnetic resonance images (MRIs) from the BraTs21[2] dataset in our experiment.
@@
Figure 7 shows an MRI of a brain with a lesion. Our objective is to segment the lesion area. Following the layout of the previous experiment, we illustrate the initial contour in (a). The initial segmentation and quasi-conformal mapping are shown in (b) and (c). From the image, it is observed that the lesion area is not connected. However, due to the topology-preserving property of the quasi-conformal mapping and the prescribed topology in the initial contour, the segmentation result is a simply-connected region. By applying two negative clicks and using the computed click map (d), we achieve the final segmentation result in (e) and its associated quasi-conformal mapping in (f).
Another example in Figure 8 depicts a more complex scenario with multiple holes inside the lesion regions. Those abnormal voids inside the area add to the segmentation complexity, often causing other methods to fail to include them. However, our method demonstrates robustness. By appropriately defining the initial contour and utilizing the topology-preserving features of the quasi-conformal mapping, our method effectively encloses a simply-connected lesion region that includes these abnormal areas, as shown in (e) and its associated mapping in (f).
@@
Figure 9 and Figure 10 show the even more challenging cases, which is to segment the lesion regions under brain MRIs with heavy noise. Following the same procedures, in this task, our methods successfully enclose the lesion region with user guidance. The result region is continuous and free of any outliers, corresponding to the prescribed topology from the initial contour input by the user. The user input clicks are effectively incorporated into the model to provide guidance for segmenting the target region. The deformation mappings are all bijective, revealing the topology-preserving feature of our interactive segmentation model.
@@ 6.3 Comprehensive analysis of QIS model
Comparison with interactive Chan-Vese segmentation model
@@
The QIS model is a deformation-based segmentation model incorporating the quasiconformal segmentation model with the click map. One advantage of the quasiconformal segmentation model is that it ensures the segmentation mask maintains the prescribed topology, avoiding topological errors and enhancing the accuracy and efficiency of the interactive segmentation model. To investigate the importance of incorporating the quasiconformal segmentation model, we compare our method with an alternative method that integrates the level-set segmentation model with the click map. Instead of utilizing the quasiconformal segmentation model, we can run the Chan-Vese segmentation model with $I^{n-1}+r^{n}M_{clicks}^{n}$ as the input image using a suitable parameter $r^{n}$, which we will refer to as interactive Chan-Vese model here after. The segmentation results of several natural images using the interactive Chan-Vese model are presented in Figure 11. The first row of Figure 11 illustrates the initial segmentation outcomes of four natural images using the conventional Chan-Vese segmentation model. The two images on the left are clean, while the two on the right are noisy. The second row displays the initial segmentation results of the same four natural images using the quasiconformal segmentation model. In both methods, the initial segmentation masks include some undesired regions and exclude certain regions of interest. Some positive clicks (in blue) and negative clicks (in red) are introduced to guide the segmentation. The third row shows the segmentation results obtained by the interactive Chan-Vese segmentation method, incorporating click maps. The segmentation masks exhibit improved accuracy in segmenting the regions of interest, though they are not entirely precise. Finally, the last row demonstrates the segmentation results obtained by the QIS model, which accurately segments the regions of interest. The topology of the segmentation mask is also consistent with that of the initial contour. Note that the clicks inputted by users are the same for both the interactive Chan-Vese segmentation model and the QIS model.
@@
Figure 12 presents another comparison using brain MRIs. The first row of Figure 12 shows the initial segmentation results of four brain MRIs using the conventional Chan-Vese segmentation model. Again, the two images on the left are clean, while the two on the right are noisy. Evidently, the initial segmentation masks contain some undesired regions and exclude some regions of interest. The second row shows the segmentation results obtained by the interactive Chan-Vese segmentation method, incorporating click maps. The clicks inputted by the users are the same as those used by the QIS model, as shown in Figure 7, Figure 8, Figure 9 and Figure 10. The segmentation masks can better segment the lesion, although they are still not accurate. The last row shows the segmentation results obtained by the QIS model, which accurately segments the lesions. The topology of the segmentation mask is consistent with that of the initial contour.
@@
Figure 11: Segmentation results on natural images obtained using the QIS model and the interactive Chan-Vese segmentation model with the integration of click maps. The left side displays the segmentation results on two clean images, while the right side shows the segmentation results on two noisy images.
@@
Figure 12: Segmentation results on brain MRIs obtained using the QIS model and the interactive Chan-Vese segmentation model with the integration of the click map. The left side displays the segmentation results on two clean brain MRIs, while the right side shows the segmentation results on two noisy brain MRIs.
Comparison with GrabCut [38]
@@
We compare the QIS model with a popular interactive segmentation model called GrabCut [38]. Similar to the QIS model, GrabCut is an interactive image segmentation algorithm that uses iterative graph cuts and Gaussian Mixture Models (GMMs) to separate foreground objects from the background. The user provides an initial bounding box around the object of interest, and the algorithm alternates between refining the segmentation using graph cuts and updating the color distribution models via GMMs. The comparison between the QIS and GrabCut methods on a noisy image capturing flowers is shown in Figure 13. In Figure 13(a), the iterative process of the QIS model is displayed. The negative clicks in each iteration, represented by red brushes, are also shown. The QIS effectively segments the noisy image in four steps with only a few brushes introduced. Figure 13(b) illustrates the iterative process of GrabCut, where the brushes introduced by the user (red for foreground and blue for background) are also shown. GrabCut produces a reasonably accurate segmentation mask after seven steps, although it is evident that it is comparatively less accurate than QIS. Figure 14 shows a further comparison between GrabCut and QIS using another noisy image capturing a road scene, with the objective of segmenting the car. The iterative process of QIS is shown in (a), where the positive and negative clicks are denoted by red and blue brushes, respectively. QIS accurately segments the car in just five steps. (b) displays the iterative process using GrabCut, with the segmentation result after eight steps exhibited, which is comparatively less accurate than QIS.
@@
Figure 14: Another comparison between the QIS and GrabCut segmentation methods. The top demonstrates the iterative segmentation process using the QIS method, with positive and negative clicks labeled as blue and red brushes, respectively. The final segmentation result is achieved in 5 steps. The bottom depicts the iterative process using the GrabCut method, with positive and negative brushes labeled as blue and red, respectively. The segmentation results after 8 steps are displayed.
@@ 7 Conclusion
In this paper, we propose the quasiconformal interactive segmentation (QIS) model, with the incorporation of user interactions. User interactions are defined by two actions: a positive click and a negative click, represented by left and right clicks, respectively. These clicks serve to indicate the regions that the user wants to include or exclude from the segmentation process. Clicks are transformed into a click map, which is then integrated into the proposed model using a novel interactive segmentation energy term that we have designed. The segmentation mask is obtained by deforming a template mask with the same topology as the object of interest using an orientation-preserving quasiconformal mapping. This approach helps to avoid topological errors in the segmentation results. We provide a thorough analysis of the proposed model, including theoretical support for the ability of QIS to include or exclude regions of interest or disinterest based on the user’s indication. To evaluate the performance of QIS, we conduct experiments on synthesized images, medical images, natural images, and noisy natural images. Regarding the computational efficiency, our method is not the fastest as presented in Appendix D. While our method may not be the fastest, as noted in Appendix D, it offers superior segmentation quality, representing a reasonable trade-off. Efforts are ongoing to further enhance its speed and efficiency.